OpenStudy (anonymous):

implicit differentiation: (e^y)(cosx)=1+sin(xy) derivative of both sides: (e^y (-sinx)+cosx(e^y))=sinxcosy+sinycosx now what?

OpenStudy (anonymous):

Are you trying to get dy/dx?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Then y' needs to be showing up in some of those functions via chain rule.

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

If you differentiate y with respect to x, you assume y is a function of x, so you use chain rule and multiply by y'

OpenStudy (anonymous):

e.g. \[\large \frac{d}{dx} e^y = y'e^y\]

OpenStudy (anonymous):

i have no idea man lol

OpenStudy (anonymous):

Also, sin(xy) is not the same as sin(x)sin(y), so you can't use the product rule like that. You will have to use product rule while using chain rule on the xy in there. It does get pretty messy...

OpenStudy (anonymous):

oooh i didnt know that

OpenStudy (anonymous):

Are you in Calc-I or Calc-II? This is a pretty nasty one for a Calc-I course I would think.

OpenStudy (anonymous):

calc 1

OpenStudy (anonymous):

Yikes. Your professor is brutal. Ok, let's step through this a little at a time. Can you take the derivative of the left side using what I told you about using chain rule on y-as-a-function-of-x?

OpenStudy (anonymous):

so well have (e^y) d/dx(cos) + (cosx) d/dx(e^y)

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

chain rule for e^y equals 0(e)^(y-1) ??

OpenStudy (anonymous):

No, d/dx of e^y = y'e^y.

OpenStudy (anonymous):

It's the derivative of e^y, which is e^y, times the derivative of y, which is y'.

OpenStudy (anonymous):

For chain rule: when you have a function-of-a-function, you multiply the derivatives of both functions together.

OpenStudy (anonymous):

i didnt know that

hartnn (hartnn):

d/dx of e^y = y'e^y. applies chain rule...

OpenStudy (anonymous):

so well have e^y(-sinx)+cosx(y'e^y)

OpenStudy (anonymous):

Good, now for the right side, remember to use product rule on the xy, while doing chain rule on sin(xy).

OpenStudy (anonymous):

(sinx) d/dx (siny)+ (siny) d/dx( sinx)

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

tsk tsk, can't do that.

OpenStudy (anonymous):

nm that well have d/dx(1) + sin [x d/dx (y) + y d/dx (x)]

OpenStudy (anonymous):

For chain rule, multiply derivative of sin by the derivative of xy, for the derivative of xy, use product rule.

OpenStudy (anonymous):

That looks good, but don't forget to d/dx the sin itself too.

OpenStudy (anonymous):

ya just noticed that

OpenStudy (anonymous):

so we should have (cosx)[(x)dy/dx + (y)(1)]

OpenStudy (anonymous):

Ok, to summarize, you should have this by this point: \[\large e^ycos(x)=1+sin(xy)\] \[\large \frac{d}{dx} e^ycos(x) = e^y(-sin(x))+cos(x)y'e^y\] \[\large \frac{d}{dx}( 1+sin(xy)) = 0+cos(xy)(xy'+y)\]

OpenStudy (anonymous):

Cos(xy), not just cos(x)

OpenStudy (anonymous):

Now set the left-side derivative equal to the right-side derivative and simplify.

OpenStudy (anonymous):

e^y(-sinx)+cosx(y'e^y)=(cosxy) (x)dy/dx + (y) -y e^y(-sinx)+cosx(y'e^y) - y = cos(xy)dy/dx divide cosxy (e^y(-sinx)+cosx(y'e^y) - y)/cos(xy) = dy/dx

OpenStudy (anonymous):

Ermmm, I don't think that works because you still have y' mixed up with the rest of the junk.

OpenStudy (anonymous):

i want this problem to end already lol

OpenStudy (anonymous):

Yeah, it's a beast! The equation simplifies to \[\large -(e^y)(sin(x))+(cos(x))(e^y)(y') = (x)(cos(xy))(y')+(y)cos(xy)\] Then get all the groups of y' on one side and everything else on the other \[\large (cos(x))(e^y)(y') - (x)(cos(xy))(y') = (y)cos(xy) + (e^y)(sin(x))\]

OpenStudy (anonymous):

i see then we factor maybe?

OpenStudy (anonymous):

Yes, you have to factor out y' so you can solve for it \[\large [(cos(x))(e^y) - (x)(cos(xy))](y') = (y)cos(xy) + (e^y)(sin(x))\]

OpenStudy (anonymous):

One last step to isolate the y' . . \[\large y' = \frac{ycos(xy) + e^ysin(x)}{e^y cos(x) - xcos(xy)}\]

OpenStudy (anonymous):

CliffSedge your a lifesaver thanks for showing me the steps

OpenStudy (anonymous):

You're welcome. Tell your professor I called him, "brutal." This was some serious differentiation. It should all be things you know how to do, but it's a damn lot of it all together! Whew!